# Li Tana,b , Shengrong Wanga,c

### v1

@@ 4 Propagation of Chaos
Since the coefficients of (2.5) depend on the distributions of state variables, we use stochastic interacting particle systems to approximate it. Firstly, considering a non-interacting particle system is given by
@@
$\mu^{X,N}_{t}(\cdot)=\frac{1}{N}\sum_{j=1}^{N}\delta_{X_{t}^{j,N}}(\cdot),$
@@
[17, Lemma 3.2] For two empirical measures $\mu_{t}^{N}=\frac{1}{N}\sum\limits_{j=1}\limits^{N}\delta_{Z^{1,j,N}_{t}}$ and $\nu_{t}^{N}=\frac{1}{N}\sum\limits_{j=1}\limits^{N}\delta_{Z_{t}^{2,j,N}}$ , then
$\mathbb{E}\left(\mathbb{W}_{q}(\mu_{t}^{N},\nu_{t}^{N})\right)\leq\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\bigg{\lvert}Z_{t}^{1,j,N}-Z_{t}^{2,j,N}\bigg{\rvert}^{q}\bigg{)}^{1/q}.$
@@
Let Assumptions 2.2-2.5 hold. For $\overline{p}>1/H$, we have
@@
where $C$ is a positive constant depending on ${\overline{p}},T,L,H,\xi$.
@@
We will proceed in the same way in Lemma 3.1. By Assumption 2.4 and Lemma 2.7, we get
$|X_{t}^{i,N}|^{\overline{p}}=|D(X_{t-\tau}^{i,N})+X_{t}^{n}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\leq\lambda|X_{t-\tau}^{i,N}|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}},$
for any $t\in[0,T]$. We immediately get
$\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\parallel\xi\parallel^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}}}\sup_{0\leq r\leq t}|X_{r}^{i,N}-D(X_{r-\tau}^{i,N})|^{\overline{p}}$
By the Itô formula, we get
@@
$+p\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\big{\langle}X_{s}^{i,N}-D(X_{s-\tau}),b\big{(}X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N}\big{)}\big{\rangle}{\rm d}s$
$+{\overline{p}}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{p-1}\sigma\big{(}\mu_{s}^{X,N}\big{)}{\rm d}B_{s}^{H}$
$+{\overline{p}}({\overline{p}}-1)\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\sigma\big{(}\mu_{s}^{X,N}\big{)}D_{s}^{\phi}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|{\rm d}s$
=:absent:=:= : $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+I_{1}(t)+I_{2}(t)+I_{3}(t).$
Then,
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\bigg{)}\leq\mathbb{E}|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{1}+I_{2}+I_{3}\bigg{)}$
By Assumptions 2.3, 2.4, and the Young’s inequality, we get
@@
$\leq$ $C{\overline{p}}\mathbb{E}\int_{0}^{t}[|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{p-2}\times(1+|X_{s}^{i,N}|^{2}+|X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})))^{2}]{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}+(1+|X_{s}^{i,N}|^{2}+|X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0}))^{2})^{{\overline{p}}/2}]{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t\wedge\nu_{R}}[(1+|X_{s}^{i,N}|^{\overline{p}}+|X_{s-\tau}^{i,N}|^{\overline{p}}+\big{(}\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})\big{)}^{{\overline{p}}}]{\rm d}s.$
By the Hölder inequality, Young’s inequality and Assumption 2.5 and in the way in (3.3), we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{2}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-1}|\sigma(\mu_{s}^{X,N})|{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}((\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0}))^{\overline{q}}+1){\rm d}s.$
By Assumption 2.5, Remark 2.1, the Young’s inequality, we get
@@
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}\bigg{[}|\sigma\big{(}\mu_{s}^{X,N}\big{)}|^{2}\int_{0}^{s}H(2H-1)|r-s|^{2H-2}{\rm d}r\bigg{]}^{{\overline{p}}/2}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+2CHT^{2H-1}\mathbb{E}\int_{0}^{t}\lvert\sigma\big{(}\mu_{s}^{X,N}\big{)}\rvert^{\overline{p}}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[(|X_{s}^{n}|^{\overline{p}}+|X_{s-\tau}^{n}|^{\overline{p}}+\big{(}\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})\big{)}^{\overline{p}}]{\rm d}s$
$=$ $C\mathbb{E}\int_{0}^{t}[(|X_{s}^{n}|^{\overline{p}}+|X_{s-\tau}^{n}|^{\overline{p}}+C\mathbb{E}\int_{0}^{t}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{r}^{j,N}|^{q}\bigg{)}^{{\overline{p}}/q}{\rm d}r,$
where for the last part, we uses the fact that
$\mathbb{W}_{\theta}(\mu_{s}^{X,N},\delta_{0})=\mathcal{W}_{q}(\mu_{s}^{X,N})=\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j,N}|^{q}\bigg{)}^{1/q}.$
Since all $j$ are identically distributed, by the Minkowski inequality and Lemma 4.1, we get
$\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j,N}|^{q}\bigg{)}^{{\overline{p}}/q}\leq\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\parallel|X_{s}^{j,N}|^{q}\parallel_{L^{1/q}}\bigg{)}^{\overline{p}/q}=\bigg{(}\frac{1}{N}\sum_{j=1}^{N}(\mathbb{E}|X_{s}^{j,N}|)^{q}\bigg{)}^{\overline{p}/q}=\mathbb{E}|X_{s}^{i,N}|^{\overline{p}}.$
@@
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s,$
Thus, we get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s.$
For $t\in[0,\tau]$, we immediately get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\bigg{)}\leq C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s\wedge\nu_{R}}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s,$
since $\xi\in L_{\mathcal{F}_{0}}^{\overline{p}}([-\tau,0];\mathbb{R}^{d})$. We get the desired result by Gronwall inequality. By induction, we can derive the assertion. ∎
@@
Let Assumption 2.4 hold. For $\overline{p}>1/H$ and $\overline{p}>2p$, we have
@@
where $C$ is a positive constant depending on $p,T,H,L,q$ but independent of $N$.
@@
$\Theta_{t}^{i}=X_{t}^{i}-D(X_{t-\tau}^{i})-X_{t}^{i,N}+D(X_{t-\tau}^{i,N}).$
By Itô formula, we get
$|\Theta_{t}^{i}|^{p}-|\Theta_{0}^{i}|^{p}$
$\leq$ $p\int_{0}^{t}|\Theta_{s}^{i}|^{p-1}\bigg{[}b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\bigg{]}{\rm d}s$
$+p\int_{0}^{t}|\Theta_{t}^{i}|^{p-1}\bigg{[}\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\bigg{]}{\rm d}B_{s}^{H,i}$
$+p(p-1)\int_{0}^{t}|\Theta_{s}^{i}|^{p-2}\bigg{(}\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\bigg{)}\bigg{[}\int_{0}^{s}\phi(s,r)(\sigma(\mathcal{L}_{X_{r}^{i}})-\sigma(\mu_{r}^{X,N})){\rm d}r\bigg{]}{\rm d}s$
=:absent:=:= : $Q_{1}(t)+Q_{2}(t)+Q_{3}(t).$
For any $t\in[0,T]$, by Assumption 2.4, the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{1}(s)\bigg{)}$
$=$ $p\int_{0}^{t}|\Theta_{s}^{i}|^{p-1}\bigg{[}b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\bigg{]}{\rm d}s$
$\leq$ $p\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{p-2}\langle\Theta_{s}^{i},b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
$\leq$ $C\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}[|X_{s}^{i}-X_{s}^{i,N}|^{p}+|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}+(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}]{\rm d}s.$
In the same way in (3.3), by Assumption 2.5, the Young’s inequality and the Hölder inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{2}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{\overline{p}-1}|\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})|{\rm d}s$
$\leq$ $C\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s.$
By Assumption 2.5, Remark 2.1, the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{3}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+2CHT^{2H-1}\mathbb{E}\int_{0}^{t}\lvert\sigma\big{(}\mathcal{L}_{X_{u}^{i}}\big{)}-\sigma\big{(}\mu_{u}^{X,N}\big{)}\rvert^{p}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[(|\Theta_{s}^{i}|^{p}+\big{(}\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})\big{)}^{p}]{\rm d}s.$

### v2

@@ 4 Propagation of Chaos
Since the coefficients of (2.5) depend on the distributions of state variables, we use stochastic interacting particle systems to approximate it. Firstly, considering a non-interacting particle system given by
@@
$\mu^{X,N}_{t}=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}},$
@@
[16, Lemma 3.2] For two empirical measures $\mu_{t}^{N}=\frac{1}{N}\sum\limits_{i=1}\limits^{N}\delta_{Z^{1,i,N}_{t}}$ and $\nu_{t}^{N}=\frac{1}{N}\sum\limits_{i=1}\limits^{N}\delta_{Z_{t}^{2,i,N}}$, then
$\mathbb{E}\mathbb{W}_{2}(\mu_{t}^{N},\nu_{t}^{N})\leq\mathbb{E}\bigg(\frac{1}{N}\sum_{i=1}^{N}\bigg\lvert Z_{t}^{1,i,N}-Z_{t}^{2,i,N}\bigg\rvert^{2}\bigg)^{1/2}.$
@@
Let Assumptions 2.2-2.5 hold. For $\overline{p}\geq 2$, we have
@@
where $C$ is a positive constant depending on ${\overline{p}},T,L,H,\xi$ but independent of $N$.
@@
We will proceed the proof in the similar way as Theorem 3.1. Firstly, Assumption 2.3 and Lemma 2.7 imply that
(4.3) $\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\parallel\xi\parallel^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}}}\sup_{0\leq s\leq t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}.$
We now concentrate on the estimation of the second term. The application of the fractional Itô formula leads to
@@
$+\overline{p}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\big\langle X_{s}^{i,N}-D(X_{s-\tau}),b\big(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N}\big)\big\rangle{\rm d}s$
$+{\overline{p}}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}-1}\|\sigma\big(\mu_{s}^{X,N}\big)\|{\rm d}B_{s}^{H}$
$+{\overline{p}}({\overline{p}}-1)H(2H-1)\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r\bigg){\rm d}s$
$=:$ $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+I_{1}(t)+I_{2}(t)+I_{3}(t).$
By means of Assumptions 2.3-2.4 and the Young’s inequality, we conclude
@@
$\leq$ $C\mathbb{E}\int_{0}^{t}[|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}+(1+|X_{s}^{i,N}|^{2}+|X_{s-\tau}^{i,N}|^{2}+\mathbb{W}_{2}^{2}(\mu_{s}^{X,N},\delta_{0}))^{{\overline{p}}/2}]{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[(1+|X_{s}^{i,N}|^{\overline{p}}+|X_{s-\tau}^{i,N}|^{\overline{p}}+\mathbb{W}_{2}^{\overline{p}}(\mu_{s}^{X,N},\delta_{0})]{\rm d}s.$
Utilizing Assumption 2.5, the Hölder inequality, the Young’s inequality and the similar way as achieving (3.6), it produces
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}I_{2}(s)\bigg)=$ ${\overline{p}}\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}-1}\|\sigma\big(\mu_{s}^{X,N}\big)\|{\rm d}B_{s}^{H}$
$\leq$ $C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{{\overline{p}}}\right)+C_{\epsilon}\mathbb{E}\int_{0}^{t}|X_{s-\tau}^{i,N}|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}\mathbb{W}_{2}^{\overline{p}}(\mu_{s}^{X,N},\delta_{0}){\rm d}s+C,$
where $0<\epsilon<1$. Again by Assumption 2.5, the Young’s inequality and the Fubini theorem, it results in
@@
$=$ ${\overline{p}}({\overline{p}}-1)H(2H-1)\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r\bigg){\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}|X_{s-\tau}^{i,N}|^{\overline{p}}{\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\mathbb{E}|X_{s}^{i,N}|^{\overline{p}}{\rm d}s,$
where we have used the fact that
$\mathbb{W}_{2}(\mu_{s}^{X,N},\delta_{0})=\mathcal{W}_{2}(\mu_{s}^{X,N})=\bigg(\frac{1}{N}\sum_{i=1}^{N}|X_{s}^{i,N}|^{2}\bigg)^{1/2}.$
Since all $X_{s}^{i,N}$ are identically distributed, the Minkowski inequality and Lemma 4.1 lead to
$\mathbb{E}\bigg(\frac{1}{N}\sum_{i=1}^{N}|X_{s}^{i,N}|^{2}\bigg)^{{\overline{p}}/2}\leq\bigg(\frac{1}{N}\sum_{i=1}^{N}\parallel|X_{s}^{i,N}|^{2}\parallel_{L^{\overline{p}/2}}\bigg)^{\overline{p}/2}=\bigg(\frac{1}{N}\bigg(\sum_{i=1}^{N}(\mathbb{E}|X_{s}^{i,N}|^{\overline{p}})^{2/\overline{p}}\bigg)^{\overline{p}/2}=\mathbb{E}|X_{s}^{i,N}|^{\overline{p}}.$
@@
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}\bigg)\leq$ $C+C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{{\overline{p}}}\right)+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i,N}|^{\overline{p}}\bigg){\rm d}r$
$+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i,N}|^{\overline{p}}\bigg){\rm d}r.$
Consequently, choosing some $\epsilon$ small enough to ensure $C\epsilon<\frac{1}{2}$ and using (4.3) to get
(4.4) $\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{\overline{p}}\bigg)\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i,N}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i,N}|^{\overline{p}}\bigg){\rm d}r.$
Since the singular kernel $(t-s)^{2H-1}$ is integrable on the finite interval $[0,T]$, we can bound
$\int_{0}^{t}(t-s)^{2H-1}{\rm d}s\leq C.$
Rewriting (4.4) by ignoring the regular kernel as
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{\overline{p}}\bigg)\leq C+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r}|X_{s}^{i,N}|^{\overline{p}}\bigg){\rm d}r,$
then, Lemma 2.8 implies
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{i,N}|^{\overline{p}}\bigg)\leq CE_{2H,1}((\Gamma(2H))^{1/2H}t).$
The Mittag-Leffler function, defined by the power series $E_{2H,1}((\Gamma(2H))^{1/2H}t)$, is an entire function. It is a well-established result that this series converges uniformly on any compact subset of the complex plane, and therefore on any finite interval $[0,T]$ (see, e.g.,[34]), which means $\mathbb{E}\bigg(\sup_{t\in[-\tau,T]}|X_{t}^{i,N}|^{\overline{p}}\bigg)$ is bounded. Finally, the boundness of $\mathbb{E}\bigg(\sup_{t\in[-\tau,T]}|X_{t}^{i}|^{\overline{p}}\bigg)$ can be shown similarly. This completes the proof. ∎
@@
Let Assumptions 2.2-2.5 hold. Fix any $\overline{p}>4$ in Lemma 4.2. Then, for $p\in[2,\frac{\overline{p}}{2})$, we have
@@
where $C$ is a positive constant depending on $p,T,H,L$ but independent of $N$ and $d$.
@@
$\Upsilon_{t}^{i}=X_{t}^{i}-D(X_{t-\tau}^{i})-X_{t}^{i,N}+D(X_{t-\tau}^{i,N}).$
By using the fractional Itô formula to $|\Upsilon_{t}^{i}|^{p}$, we can show
$|\Upsilon_{t}^{i}|^{p}-|\Upsilon_{0}^{i}|^{p}$
$\leq$ $p\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\langle\Upsilon_{s}^{i},b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
$+p\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\langle\Upsilon_{s}^{i},\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}$
$+p(p-1)H(2H-1)\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mathcal{L}_{X_{r}^{i}})-\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r\bigg){\rm d}s$
$=:$ $H_{1}(t)+H_{2}(t)+H_{3}(t).$
Assumptions 2.3, 2.4 and the Young’s inequality lead to
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}H_{1}(s)\bigg)=p\mathbb{E}\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\langle\Upsilon_{s}^{i},b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{t}^{i}-X_{t}^{i,N}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}){\rm d}s.$
As the same way of achieving (3.6), we use Assumption 2.5, the Young’s inequality and the Hölder inequality to get
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}H_{2}(s)\bigg)=$ $p\mathbb{E}\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\langle\Upsilon_{s}^{i},\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}$
$\leq$ $C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t}|\Upsilon_{s}^{i}|^{{p}}\right)+C\mathbb{E}\int_{0}^{t}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}){\rm d}s.$
Further, Assumptions 2.3, 2.5, Remark 2.1, the Young’s inequality and Fubini theorem give
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}H_{3}(s)\bigg)$
$=$ $p(p-1)H(2H-1)\int_{0}^{t}|\Upsilon_{s}^{i}|^{p-2}\bigg(\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mathcal{L}_{X_{r}^{i}})-\sigma(\mu_{r}^{X,N})\|^{2}{\rm d}r\bigg){\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|\Upsilon_{s}^{i}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(t-s)^{2H-1}\mathbb{W}_{2}^{p}\big(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}\big){\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{t}^{i}-X_{t}^{i,N}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(t-s)^{2H-1}\mathbb{W}_{2}^{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}){\rm d}s.$